For the claim, "Colorblindness occurs in 9% of men," find the p-value for the test statistic z=2.05.
At α=0.1 significance level, would we reject H0?

Answer by rothauserc(4718)

Ho: x = 9%
H1: x not = 9%
:
this is a two-tailed test
:
p-value = (1 - 0.9798) * 2 = (0.0202) * 2 = 0.0404
:
since our p-value < 0.10, we reject Ho
: